### v1

@@ 1 Introduction
This paper introduces an incremental identification approach specifically designed for general dynamic hybrid models. To address a deficiency of previous approaches, we introduce a regularization term in the dynamic parameter estimation problem to prevent large discontinuities in the flux profiles determined at the first stage of the incremental identification. Such large discontinuities have no physical basis, and, consequently, affect the quality of the data-driven elements identified at the second stage of the incremental identification. This problem is particularly pronounced when the amount of available experimental data is limited, which is often the case in practical applications. Furthermore, we consider generally unkown model parts with or without physical meaning.
We apply the new incremental identification approach in three case studies. The first case study illustrates the identification approach by focusing on a chemical reactor. The second case study focuses on dynamic bioreactor models and illustrates the development of a dynamic hybrid model with a limited amount of data. The third case study focuses on a real research plant. In this case study, we identify a dynamic hybrid model which we embed as a controller model in a MPC framework to control the research plant.
@@ 2 Dynamic Hybrid Modeling Approach
The model in Equation (1) is under-specified due to the inclusion of an additional set of unknown variables, $\boldsymbol{p}(t)\in\mathbb{R}^{N_{p}}$, which are difficult—or even impossible—to describe mechanistically. These variables may represent, for example, the rates of chemical reactions or fluxes associated with mass and heat transfer phenomena, particularly in cases where the underlying physics is not fully understood. However, they need not have a direct physical interpretation. In some cases, they may serve as correction or “error” terms applied to specific mechanistic expressions or even to entire equations.
@@ 2.1 Regularized Dynamic Parameter Estimation
$\boldsymbol{0}$ $\geq\boldsymbol{c}(\boldsymbol{x}(t),\boldsymbol{y}(t),\boldsymbol{u}(t),\boldsymbol{p}(t)),t\in\mathcal{T},$ (2f)
with the time horizon $\mathcal{T}=[t^{\mathrm{meas}}_{1},t^{\mathrm{meas}}_{N_{\mathrm{meas}}}]$, the initial states $\boldsymbol{x}_{0}\in\mathbb{R}^{N_{x}}$, the output variables $\boldsymbol{z}_{j}\in\mathbb{R}^{N_{z}}$ at the distinct time points $t_{j}\in\mathcal{T}^{\mathrm{meas}}$, the output equations $\boldsymbol{h}:\mathcal{X}\rightarrow\mathbb{R}^{N_{z}}$, path and terminal (equality or inequality) constraints $\boldsymbol{c}:\mathcal{X}\rightarrow\mathbb{R}^{N_{g}}$, and the regularization function $R:\mathbb{R}^{N_{p}}\rightarrow\mathbb{R}$. $\boldsymbol{W}_{j}^{z}\in\mathbb{R}^{N_{z}\times N_{z}}$, $j\in\{1,...,N_{\mathrm{meas}}\}$ is the covariance matrix of the measurements $\tilde{\boldsymbol{z}}_{j}$ at time point $t_{j}$.
@@
The dynamic optimization problem (2) may be ill-behaved, with multiple time profiles $\boldsymbol{p}(t)$ resulting in the same or very similar fits of predictions to experimental data. This is especially likely when only limited experimental data is available. In practice, this situation often results in the optimal solution exhibiting large jumps in the values of $\boldsymbol{p}(t)$ from one discretization interval to the next. From the point of view of the hybrid model identification, this is problematic: as we shall see in Section 2.3, the parameters $\boldsymbol{p}(t)$ will ultimately be expressed, e.g., via machine learning models, as continuous functions (e.g. ANNs) of subsets of the model variables $\boldsymbol{x}(t)$, $\boldsymbol{y}(t)$, $\boldsymbol{u}(t)$ which typically represent physical quantities that may not exhibit such discontinuous behavior. Overall, the discrepancy between the smoothness of $\boldsymbol{p}(t)$ and that of $\boldsymbol{x}(t)$, $\boldsymbol{y}(t)$, $\boldsymbol{u}(t)$ may significantly affect the quality of the data-driven model that can be identified based on the results of the dynamic optimization problem.
In view of the above, the second part of the objective function (2a) is a weighted regularization term that penalizes changes in the values of the parameters $\boldsymbol{p}(t)$ between successive discretization intervals:
@@
The values of the regularization parameters $\boldsymbol{W}_{k}^{R}$ need to be selected carefully to avoid over-penalizing changes in $\boldsymbol{p}(t)$ in time periods where there may, indeed, occur variations in the state variables $\boldsymbol{x}(t)$, $\boldsymbol{y}(t)$, e.g. caused by large discontinuous changes in the values of the MVs $\boldsymbol{u}(t)$.
@@ 2.4 Hybrid Model Integration
As a final check on the quality of the identified hybrid model (4), we can use it to simulate each of the original experimental datasets, and compute the corresponding value of the weighted deviation between model prediction and experimental measurements. We note that these deviations may be different (not necessarily worse) than those originally determined via the optimal solutions of the dynamic optimization problem (2). Such differences may arise because: (1) the data-driven model derived at Step 3 (cf. Section 2.3) may not be a sufficiently accurate representation of the dependence of $p$ on the other system variables and (2) at Step 1 (cf. Section 2.1), the parameter profiles $\boldsymbol{p}(t)$ were approximated as a piecewise constant functions of time, but now they are treated as continuous functions of the other system variables. In principle, a fine-tuning of the hybrid model could be performed at this stage by solving a dynamic optimization problem incorporating multiple instances of the hybrid model (4), one for each experimental dataset. The parameters of the data-driven model (e.g. ANN weights and biases) would be treated as optimization decision variables. However, the solution of such a problem is likely to be very computationally expensive for anything other than simple models.
The major advantages of this the methodology described in this Section 2 result from the identification being carried out in two stages. Step 1 (cf. Section 2.1) determines time profiles for those quantities $\boldsymbol{p}(t)$ that are difficult to characterize mechanistically. Step 3 (cf. Section 2.3) then establishes the relation between these quantities and the other system variables. This separation allows each step to be evaluated independently. If the measured and estimated states from Step 1 diverge significantly, the model structure from Step 1 must be revised as any such mismatch is independent of the data-driven structure used later. At Step 3, the data-driven model is determined entirely from a data table established at Step 1 without the need to take any further account of the original mechanistic model (1). This greatly facilitates hyper-parameter optimization and training using standard libraries and approaches like TensorFlow, something which is problematic for any simultaneous identification approach. In any case, if the data-driven model is unable to produce a satisfactory prediction of the parameters $\boldsymbol{p^{*}}(t)$, its model structure should be revised.
@@ 3 Case Studies
  • Case Study 3: Research Plant with MPC — This real-world application demonstrates how the identified hybrid model can be embedded within a MPC framework for effective process control.
@@ 3.2 Illustrative Chemical Reactor Case Study
Fig. 2: Results of dynamic parameter estimation problem (Step 1) for chemical reactor case study. blue dots: Noisy measurement data. Red, solid lines: optimal profiles, i.e., solution of the parameter optimization problem. Black, dashed lines: profiles of mechanistic model for parameters $p_{1}$, $p_{2}$, $p_{3}$. (a) Height profile and data. (b) Parameter $p_{1}$ profile and data. (c) Concentration profile and data. (d) Parameter $p_{2}$ profile and data. (e) Temperature profile and data. (f) Parameter $p_{3}$ profile and data.
@@
$\Delta H$ $-5\times 10^{4}$ kJ/kmol
@@ 3.2.1 Pseudo-Experimental Data Generation
To emulate real-life data, the mechanistic model (5) is simulated with different initial conditions for the differential states and different MV profiles (outlet flowrate $F_{\mathrm{out}}$ and cooling water temperature $T_{\mathrm{c}}$). In total, 8 sets of initial conditions and MV profiles were used, spanning 20 hours with a 1-second discretization. We add 2% Gaussian white noise to the output values (differential states). The noisy output values and the corresponding input functions represent the measurement data used for the parameter estimation in Step 1.
@@ 3.2.5 Step 4: Hybrid Model Integration
To analyze the performance of the hybrid model, we simulate the hybrid model (7) for new MV profiles and compare the simulation results with those obtained using the original mechanistic model (5), using the same initial states and MV profiles. Figure 4 shows the corresponding simulation results. The profiles of $p_{2}$ and $p_{3}$ as predicted by the data-driven models (ANNs) are shown in Figures 4(b) and 4(d). We observe that the hybrid model simulation results closely match the simulation results of the original model, even with new MV profiles. The differential states show divergence only after approximately 2 hours due to error accumulation at each integration step. As seen in Figures 4(b) and 4(d), the parameters $p_{2}$ and $p_{3}$ are accurately predicted in the hybrid simulation, even over a long time horizon. The hybrid model demonstrates good performance, effectively capturing the dynamics of the system.
@@
Fig. 4: Simulation results with hybrid model compared to simulation results with mechanistic model for chemical reactor case study. Red, solid lines: simulation results with hybrid model. Black, dashed lines: simulation results with mechanistic model. (a) Concentration profiles. (b) Temperature profiles. (c) Parameter $p_{2}$ profiles. (d) Parameter $p_{3}$ profiles.
@@ 3.3 Bioreactor Case Study
We consider a complex bioreactor system in this case study. The mechanistic model of the bioreactor is available in model library of gPROMS FormulatedProducts888gPROMS FormulatedProducts, version 2023.2.0, www.siemens.com/global/en/products/automation/industry-software/gproms-digital-process-design-and-operations/gproms-modelling-environments/gproms-formulatedproducts.html, accessed: 26.05.2025 and consists of 30 differential equations and over 2500 algebraic equations. We use implementation alternative 1 for this case study (cf. Section 3.1).
@@ 3.3.1 Pseudo-Experimental Data Generation
To emulate measurement data, we simulate the mechanistic model with different initial states for the asparagine concentration $c_{\mathrm{A0}}$ of 1, 7, and 20 mmol/L. The output data includes the glucose concentration $c_{\mathrm{G}}$, lactate concentration $c_{\mathrm{L}}$, asparagine concentration $c_{\mathrm{A}}$, and cell viability $v$, sampled daily over a 14-day period. We add 5% Gaussian white noise to the output values to obtain realistic measurement data. The specific cell growth rate $p_{1}$ and specific cell death rate $p_{2}$ kinetics are assumed to be unknown and need to be estimated.

### v2

@@ Abstract
This is the author’s accepted manuscript. The final published version can be accessed via the following DOI: https://doi.org/10.1016/j.compchemeng.2025.109413.
@@ 1 Introduction
This paper introduces an incremental identification approach specifically designed for general dynamic hybrid models. To address a deficiency of previous approaches, we introduce a regularization term in the dynamic parameter estimation problem to prevent large discontinuities in the flux profiles determined at the first stage of the incremental identification. Such large discontinuities have no physical basis, and, consequently, affect the quality of the data-driven elements identified at the second stage of the incremental identification. This problem is particularly pronounced when the amount of available experimental data is limited, which is often the case in practical applications. Furthermore, we consider generally unknown model parts with or without physical meaning.
We apply the new incremental identification approach in three case studies. The first case study illustrates the identification approach by focusing on a chemical reactor. The second case study focuses on dynamic bioreactor models and illustrates the development of a dynamic hybrid model with a limited amount of data. The third case study focuses on a real-world research plant. In this case study, we identify a dynamic hybrid model which we embed as a controller model in a MPC framework to control the research plant.
@@ 2 Dynamic Hybrid Modeling Approach
The model in Equation (1) is under-specified due to the inclusion of an additional set of unknown variables, $\boldsymbol{p}(t)\in\mathbb{R}^{N_{p}}$, which are difficult—or even impossible—to describe mechanistically. These variables may represent, for example, the rates of chemical reactions or fluxes associated with mass and heat transfer phenomena, particularly in cases where the underlying physics is not fully understood. However, they need not have a direct physical interpretation. In some cases, they may serve as correction or “error” terms applied to specific mechanistic expressions or even to entire equations, e.g., using the following form:
$\frac{d\boldsymbol{x}}{dt}(t)=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{y}(t),\boldsymbol{u}(t))+\boldsymbol{p}(t).$
Here, the parameters $\boldsymbol{p}(t)$ serve as additive corrections to account for unmodeled dynamics or uncertain phenomena. This structure is particularly useful when detailed mechanistic knowledge is unavailable or when the unknowns lack a clear physical interpretation. Our incremental identification approach is flexible and supports both embedded and additive formulations.
In large-scale systems where the unknown profiles $\boldsymbol{p}(t)$ lack a clear physical interpretation, determining which differential or algebraic equations should include these parameters is closely tied to the concept of identifiability [23]. Model structures with guaranteed system-theoretical properties can be determined using semi-infinite programming approaches, as in [24]. A systematic methodology for selecting the appropriate placement and number of unknown profiles is beyond the scope of this work. Instead, we assume that the structure of the hybrid model (1) is already determined by the model developer.
@@ 2.1 Regularized Dynamic Parameter Estimation
$\boldsymbol{0}$ $\geq\boldsymbol{c}(\boldsymbol{x}(t),\boldsymbol{y}(t),\boldsymbol{u}(t),\boldsymbol{p}(t)),\ \forall t\in\bar{\mathcal{T}},$ (2f)
with the time horizon $\mathcal{T}=[t^{\mathrm{meas}}_{1},t^{\mathrm{meas}}_{N_{\mathrm{meas}}}]$, the initial states $\boldsymbol{x}_{0}\in\mathbb{R}^{N_{x}}$, the output variables $\boldsymbol{z}_{j}\in\mathbb{R}^{N_{z}}$ at the distinct time points $t_{j}\in\mathcal{T}^{\mathrm{meas}}$, the output equations $\boldsymbol{h}:\mathcal{X}\rightarrow\mathbb{R}^{N_{z}}$, inequality constraints $\boldsymbol{c}:\mathcal{X}\rightarrow\mathbb{R}^{N_{g}}$, that need to be satisfied over a subset $\bar{\mathcal{T}}\subseteq\mathcal{T}$ of the time domain $\mathcal{T}$, and the regularization function $R:\mathbb{R}^{N_{p}}\rightarrow\mathbb{R}$. $\boldsymbol{W}_{j}^{z}\in\mathbb{R}^{N_{z}\times N_{z}}$, $j\in\{1,...,N_{\mathrm{meas}}\}$ is the covariance matrix of the measurements $\tilde{\boldsymbol{z}}_{j}$ at time point $t_{j}$.
@@
The dynamic optimization problem (2) may be ill-behaved, with multiple time profiles $\boldsymbol{p}(t)$ resulting in the same or very similar fits of predictions to experimental data. This is especially likely when only limited experimental data is available. In practice, this situation often results in the optimal solution exhibiting large jumps in the values of $\boldsymbol{p}(t)$ from one discretization interval to the next. From the point of view of the hybrid model identification, this is problematic: as we shall see in Section 2.3, the parameters $\boldsymbol{p}(t)$ will ultimately be expressed, e.g., via machine learning models, as continuous functions (e.g. ANNs) of subsets of the model variables $\boldsymbol{x}(t)$, $\boldsymbol{y}(t)$, $\boldsymbol{u}(t)$ which typically represent physical quantities that may not exhibit such discontinuous behavior. Overall, the discrepancy between the smoothness of $\boldsymbol{p}(t)$ and that of $\boldsymbol{x}(t)$, $\boldsymbol{y}(t)$, $\boldsymbol{u}(t)$ may significantly affect the quality of the data-driven model that can be identified based on the results of the dynamic optimization problem. In view of the above, the second part of the objective function (2a) is a weighted regularization term that penalizes changes in the values of the parameters $\boldsymbol{p}(t)$ between successive discretization intervals:
@@
The values of the regularization parameters $\boldsymbol{W}_{k}^{R}$ need to be selected carefully to avoid over-penalizing changes in $\boldsymbol{p}(t)$ in time periods where there may, indeed, occur variations in the state variables $\boldsymbol{x}(t)$, $\boldsymbol{y}(t)$, e.g. caused by large discontinuous changes in the values of the MVs $\boldsymbol{u}(t)$. Section 3.2.6 provides a comparative analysis illustrating the impact of the regularization on the smoothness and quality of the estimated profiles.
@@ 2.4 Hybrid Model Integration
As a final check on the quality of the identified hybrid model (4), we can use it to simulate each of the original experimental datasets, and compute the corresponding value of the weighted deviation between model prediction and experimental measurements. We note that these deviations may be different (not necessarily worse) than those originally determined via the optimal solutions of the dynamic optimization problem (2). Such differences may arise because: The data-driven model derived at Step 3 (cf. Section 2.3) may not be a sufficiently accurate representation of the dependence of $p$ on the other system variables. Moreover, at Step 1 (cf. Section 2.1), the parameter profiles $\boldsymbol{p}(t)$ were approximated as piecewise constant functions of time, but now they are treated as continuous functions of the other system variables. In principle, a fine-tuning of the hybrid model could be performed at this stage by solving a dynamic optimization problem incorporating multiple instances of the hybrid model (4), one for each experimental dataset. The parameters of the data-driven model (e.g. ANN weights and biases) would be treated as optimization decision variables. However, the solution of such a problem is likely to be very computationally expensive for anything other than simple models.
A final confirmation of the hybrid model’s quality can be performed via standard techniques such as $k$-fold cross validation. Ultimately, the quality of the hybrid model will depend on (a) the quality of the underlying mechanistic model, (b) the way in which the data-driven elements are introduced within this mechanistic model, and (c) the quality and amount of the available experimental data.
The major advantages of the methodology described in this Section 2 result from the identification being carried out in two stages. Step 1 (cf. Section 2.1) determines time profiles for those quantities $\boldsymbol{p}(t)$ that are difficult to characterize mechanistically. Step 3 (cf. Section 2.3) then establishes the relation between these quantities and the other system variables. This separation allows each step to be evaluated independently. If the measured and estimated states from Step 1 diverge significantly, the model structure from Step 1 must be revised as any such mismatch is independent of the data-driven structure used later. At Step 3, the data-driven model is determined entirely from a data table established at Step 1 without the need to take any further account of the original mechanistic model (1). This greatly facilitates hyper-parameter optimization and training using standard libraries and approaches like TensorFlow, something which is problematic for any simultaneous identification approach. In any case, if the data-driven model is unable to produce a satisfactory prediction of the parameters $\boldsymbol{p^{*}}(t)$, its model structure should be revised.
@@ 3 Case Studies
  • Case Study 3: Research Plant with MPC — This real-world case demonstrates how the identified hybrid model can be integrated into an MPC framework for effective process control.
@@ 3.2 Illustrative Chemical Reactor Case Study
Fig. 2: Results of dynamic parameter estimation problem (Step 1) for chemical reactor case study. Blue dots: Noisy pseudo-experimental measurement data. Red, solid lines: optimal profiles, i.e., solution of the parameter optimization problem. Black, dashed lines: mechanistic model predictions for parameters $p_{1}$, $p_{2}$, $p_{3}$. For ease of presentation, each figure represents the concatenation of 8 separate experiments, each of 2.5 h duration but with different initial conditions and MV input profiles. (a) Height $h$. (b) Parameter $p_{1}$. (c) Concentration $c$. (d) Parameter $p_{2}$. (e) Temperature $T$. (f) Parameter $p_{3}$.
@@
$\Delta H$ $5\times 10^{4}$ kJ/kmol
@@ 3.2.1 Pseudo-Experimental Data Generation
To emulate real-life data, the mechanistic model (5) is simulated with different initial conditions for the differential states and with different MV profiles (outlet flowrate $F_{\mathrm{out}}$ and cooling water temperature $T_{\mathrm{c}}$). In total, 8 sets of initial conditions and MV profiles were used, spanning a total duration of 20 hours. The MVs were varied in a piecewise constant manner over a uniform time grid with a 1-minute time interval. The values in each interval were taken to be random samples from normal distributions with means of 0.1 m3/min for $F_{out}$ and 300 K for $T_{c}$, and standard deviations of 1 % for both. To create the pseudo-experimental measurements, we add 2% Gaussian white noise to the predicted values of the differential states.
@@ 3.2.5 Step 4: Hybrid Model Integration
To analyze the performance of the hybrid model, we simulate the hybrid model (7) for new MV profiles different to those used for the generation of the pseudo-experimental data (cf. Section 3.2.1) but generated using the same randomized methodology and compare the simulation results with those obtained using the original mechanistic model (5), using the same initial states and MV profiles.
Figure 4 shows a comparison of the predictions of the hybrid model against those of the mechanistic model. As can be seen in Figures 4(b) and 4(d), the time profiles of the parameters ${p}_{2}$ and ${p}_{3}$ computed by the data-driven element (neural network) within the hybrid model are reasonably close approximations to the true profiles of the corresponding terms in the mechanistic model. However, the corresponding profiles of the state profiles shown in Figures 4(a) and 4(c) exhibit larger divergences between hybrid and mechanistic model predictions. This is to be expected: a comparison of the forms of the two models (i.e. Equations (5b) and (5c) for the mechanistic model vs. Equations (7b) and (7c) for the hybrid one) indicates that the state profile errors depend on the time integrals of the errors in the corresponding parameters $\boldsymbol{p}$. Consequently, over long simulation times, errors in the hybrid model’s predictions of the states may accumulate significantly even if the parameters $\boldsymbol{p}$ are predicted well. Moreover, whether or not the magnitudes of these state errors are increasing or decreasing at a particular point in time will depend on whether the state and parameter errors have the same sign (positive or negative). This can clearly be seen from a comparison of Figures 4(c) and 4(d). Overall, in applications where the hybrid model is deployed online over extended periods of time, the use of state estimation techniques is essential for correcting the state trajectories by taking account of real-time measurements.
@@
Fig. 4: Simulation results with hybrid model compared to simulation results with mechanistic model for chemical reactor case study. Red, solid lines: simulation results with hybrid model. Black, dashed lines: simulation results with mechanistic model. (a) Concentration profiles. (b) Parameter $p_{2}$ profiles. (c) Temperature profiles. (d) Parameter $p_{3}$ profiles.
3.2.6 Influence of Regularization in Dynamic Parameter Estimation Problem (Step 1)
Fig. 5: Solution of dynamic parameter estimation (step 1) with the regularization term (blue, dash-dotted lines) and without the regularization term (brown, dashed lines) in the objective function (2a). The measurement samples used for parameter estimation are illustrated by black dots. (a) Height profile. The correct profiles (determined via simulation of the original model) are illustrated by black, solid lines. (b) Parameter $p_{1}$ profile. (c) Concentration profile. (d) Parameter $p_{2}$ profile. (e) Temperature profile. (f) Parameter $p_{3}$ profile.
To illustrate the impact of the regularization term $R(\boldsymbol{p}(t))$ on the results the dynamic parameter estimation problem (2), we consider a slightly modified version of the problem solved in Section 3.2.2. In particular, to emulate the issues arising from sparsity of experimental data, the estimation problem takes account of only a small subset of the measurement points. Moreover, in the interests of clarity of results presentation, the uniform grid used for the piecewise constant representation of the parameter functions $p(t)$ involves time intervals of duration 1 h instead of 1 minute as is the case in Section 3.2.2.
The results with and without the regularization term (cf. Equation (3)) are shown in Figure 5. It can be seen that, while both approaches yield a good fit to the sparse measurement data $\boldsymbol{\tilde{z}}_{j}$ (cf. Figures 5(b), 5(d) and 5(f)), significant differences emerge in the smoothness and physical plausibility of the resulting profiles $\boldsymbol{p}(t)$ and $\boldsymbol{x}(t)$. Without regularization, the estimated parameter trajectories exhibit strong eratic fluctuations and deviate notably from the true simulation profiles. In contrast, the inclusion of the regularization term $R(\boldsymbol{p}(t))$ (cf. Equation (3)) leads to smoother parameter and state trajectories that more closely follow the underlying simulation data. This highlights the benefit of regularization in promoting physically meaningful solutions, especially when experimental data is limited.
It is worth noting that, for the examples considered in this paper, we assumed a diagonal structure for the matrices $\boldsymbol{W}_{k}^{R}$ used in the regularization term of Equation (3), and adopted an empirical approach for their determination. In particular, starting with matrices with zero or very small elements in the diagonal, we solved the regularized dynamic parameter estimation problem (2) and observed the resulting time profiles for $\boldsymbol{p}(t)$. If one or more of the resulting $\boldsymbol{p}$ was found to exhibit erratic behavior, we increased the corresponding elements in $\boldsymbol{W}_{k}^{R}$ and repeated the solution.
@@ 3.3 Bioreactor Case Study
We consider a complex bioreactor system in this case study. The mechanistic model of the bioreactor is implemented in gPROMS FormulatedProducts888gPROMS FormulatedProducts, version 2023.2.0, www.siemens.com/global/en/products/automation/industry-software/gproms-digital-process-design-and-operations/gproms-modelling-environments/gproms-formulatedproducts.html, accessed: 26.05.2025 and consists of 30 differential and over 2500 algebraic equations. We use implementation alternative 1 for this case study (cf. Section 3.1).
As is often the case in modeling of bioreactors, the main gap in mechanistic knowledge is related to the characterization of the underlying biological phenomena. Here, guided by domain-specific knowledge, we assume that the major such uncertainty is associated with the cell growth and cell death mechanisms. We therefore treat the specific cell growth and death rates as the parameters $p_{1}$ and $p_{2}$ that need to be characterized via data-driven elements in our hybrid model.
@@ 3.3.1 Pseudo-Experimental Data Generation
To generate pseudo-experimental data, we simulate the mechanistic bioreactor model using three distinct initial conditions for the asparagine concentration $c_{\mathrm{A0}}$: 1, 7, and 20 mmol/L. The output variables include glucose concentration $c_{\mathrm{G}}$, lactate concentration $c_{\mathrm{L}}$, asparagine concentration $c_{\mathrm{A}}$, and cell viability $v$, sampled daily over a 14-day period. To emulate realistic measurement noise, we add 5% Gaussian white noise to the output values.